OpenStudy (anonymous):

Al203. If every oxygen ion (combined with an aluminum ion) has a charge of –2, what must the charge of each aluminum ion be? How can you tell this? Also, what has to happen to an uncharged aluminum atom so that it will bond with oxygen?

OpenStudy (photon336):

unless there's a charge that's visible i.e. like some kind of poly atomic ion you must do two things: for every compound the sum of the oxidation states must sum up to zero, unless it's an ion then the sum of the oxidation states must equal the charge of that particular ion. first we count how many and what type of atoms we have Al 2 atoms Oxygen 3 atoms now we know that the charge of oxygen must be -2 so we multiply this by the number of atoms that we have so that's -2*3= -6 now what number added to -6 would give us zero. -6+6 = 0 so that gives us +6 now we have two aluminum atoms; which add up to a charge of +6 so that's +3 why? because +3*2 = 6+-6 = 0 so the oxidation state of aluminum is +3 and the oxidation state of oxygen is -2